OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1). y + 3 = −5(x − 2) y − 2 = −5(x + 3) y + 3 = −one fifth(x − 2) y − 2 = −one fifth(x + 3)

OpenStudy (aroub):

slope= 1-2/ 2+3 = -1/5 y-y0=m(x-x0) y-2=-1/5 (x+3

OpenStudy (anonymous):

thank you

OpenStudy (aroub):

anytimmee!! :)